OpenStudy (anonymous):

An open rectangular box with square base is to be made from 48 ft^2 of material. What dimensions will result in a box with the largest possible volume?

OpenStudy (anonymous):

4,4,8 Let sides be x ,a,a AS a^2+4ax=48 which gives x=(48-a^2)/4a volume=a^2 x Put value of x in this we get V=(48a-a^3)/4 differentiating wrt a and equating to zerowe get 48-3a^2=0 a^2=16 a=4 hence x=8

OpenStudy (anonymous):

Square base means length = width. V = L*W*H = L^2 * H SA = 2L^2 + 4LH = 2L(L+2H) = 48 H = 48/2L -L =(48-2L^2)/2L V = L^2 * [(48-2L^2)/2L] There's your fomula for volume in terms of 1 variable, Length. Take the derivative, solve for 0 to get critical points. CHeck those points.

OpenStudy (anonymous):

If a=4, then x would be 2 wouldn't it. (48-(4)^2)/ (4(4) = 2

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

calculation error!!!!

OpenStudy (anonymous):

thanx for correcting that

OpenStudy (anonymous):

V = 24L-L^3 That's the simplified form I get for volume. V' = 24 -3L^2 24-3L^2 = 0 3(8-L^2) = 0 gives L = sqrt(8)

OpenStudy (anonymous):

Dude its an open box

OpenStudy (anonymous):

And if L = sqrt(8), then H = (48-2L^2)/2L = 32/(2sqrt8) = 16/sqrt8

OpenStudy (anonymous):

Hmph. So minor change in SA. Same method otherwise.

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

What difference does it make it is an open box?

OpenStudy (anonymous):

I guess it means SA = L^2 + 4LH instead of 2L^2 + 4LH

OpenStudy (anonymous):

open box doesnt count top one

OpenStudy (anonymous):

@SmoothMath yess

OpenStudy (anonymous):

Oh. Haha. Duh. Okay. Thank you guys.

OpenStudy (anonymous):

it was pleasure helping u

OpenStudy (anonymous):

SA = L^2 + 4LH = 48 H = (48-L^2)/(4L) V = L^2H = L^2(38-L^2)/4L = (48L-L^3)/4 = 48L/4 -(L^3)/4 V' = 12 - 4L^2 Setting equal to 0 yields 4(3-L^2) =0 L = sqrt(3)

OpenStudy (anonymous):

Oh hold on. v' = 12 - (4/3)L^2

OpenStudy (anonymous):

again incorrect

OpenStudy (anonymous):

Lols. Okay.

OpenStudy (anonymous):

v' = 12 - (3/4)L^2

OpenStudy (anonymous):

Ah. Indeed.

OpenStudy (anonymous):

:)